OpenStudy (anonymous):

can someone helpplease in 1993 the life expectancy of males in a certain country was 70.8 years in 1997 it was 73.3 years let E represent the life expectancy in year t and let t represent the number of years since 1991 The linear function that fits the data is: E(t) = --t + - round to the nearest tenth use the function to predict the life expectancy of males in 2007 E(14) = -- round to nearest tenth

OpenStudy (anonymous):

Look at 1993 70.8 = 1993 -1991 + c = 3+ c c = 67.8 therefore E(t) = year -1991 + 67.8 Now use this to check data in the year 1997 73.3 = 1997-1991 + 67.8 73.3 = 6 + 67.8 (should that read 73.8 by the way not 73.3?) Now do the last part..., E(14) =2005 -1991 + 67.8 81.8

OpenStudy (anonymous):

no it is 73.3